OpenStudy (anonymous):

solve the equation: p(p+3)=4

OpenStudy (anonymous):

bring the 4 over p(p+3)-4=0

OpenStudy (unklerhaukus):

\[p(p+3)=4\] \[p(p+3)-4=0\] \[p^2+3p-4=0\] now use the quadratic equation

OpenStudy (anonymous):

p2 + 3p - 4 = 0?

OpenStudy (unklerhaukus):

of factor in your head using the factors of -4

OpenStudy (anonymous):

so you factor out P^2+3p-4=0 @UnkleRhaukus

OpenStudy (unklerhaukus):

what do you get/?

OpenStudy (anonymous):

you make the 4 a 2^2 right?

OpenStudy (anonymous):

(p+4)(p-1) which gives you p=-4 p=1

OpenStudy (anonymous):

@UnkleRhaukus

OpenStudy (unklerhaukus):

you make the -4=-1x4